OpenStudy (anonymous):

help plz..question 1. Are 2x+4/3x+6 and 2/3 equivalent expressions? Using complete sentences explain why or why not... question 2..Using complete sentences, explain each step in simplifying the ratio and provide the simplified form.x^2+8x/x^2+10x+16

OpenStudy (campbell_st):

well the expression can be written as \[\frac{2x + 4}{3x + 6}\] when factorised you get \[\frac{2(x +2)}{3(x + 2)}\] you need to decide if there is a common factor that can be removed to simplify the expression. If there is a common factor what's left?

OpenStudy (campbell_st):

for question 2. Factorise the numerator and denominator.... there will be a common factor that can be removed,

OpenStudy (anonymous):

question 1.cancel out +2??

OpenStudy (campbell_st):

why not just cancel (x +2) its in the numerator and denominator

OpenStudy (anonymous):

question 2 will be...x^2 + 8x)/(x^2 + 10x +16) x(x + 8)/(x + 8)(x + 2) get this by factoring each expression the (x + 8)/(x + 8) = 1 so they cancel each other and the expression simplifies to: x/(x + 2) you can prove to yourself that the factors are correct by multiplying the terms again to get the original expression

OpenStudy (campbell_st):

thats correct. to do the next part I'd show \[\frac{(x + 8)}{x + 8} \times \frac{x}{x+ 2} = \frac{x^2 + x8}{x^2 + 10x + 16}\]

OpenStudy (anonymous):

k thanks